OpenStudy (anonymous):

Ratio Test problem: Summation {n=1}^infty (n/(2n+5))^n I got L = 1/(2e^2) but I don't know where/how else to check my answer. I see that it converges since 1/(2e^2) <1.. but that's it. Any help is appreciated! Thank you :)

OpenStudy (dan815):

yeah

OpenStudy (therealmeeeee):

well @dan815 lets get this young lady a A+++++++++++++++++

OpenStudy (anonymous):

\[\sum_{1}^{\infty} (\frac{ n }{ 2n+5 })^{n}\] \[L = \frac{ 1 }{ 2e^2 }\]

OpenStudy (dan815):

are you given 1/2e^2 or did u find out that it is less than that bound

OpenStudy (therealmeeeee):

I'll assume the question is 10n2−1+2n−5n−1=2n+5n+1 eliminate the denominator by multiplying every term by (n-1)(n+1) 10+(2n−5)(n+1)=(2n+5)(n−1) simplify the expression 10+2n2−3n−5=2n2+3n−5 collect like terms gives 10−6n=0 so 6n = 10 n = 5/3

OpenStudy (anonymous):

I got 1/2e^2 and, give me a min to read that last post, haha

OpenStudy (dan815):

ratio test this converges

OpenStudy (therealmeeeee):

wait wat wat i did did not help u at all X_X

OpenStudy (anonymous):

so i went on wolframalpha to double check my answer, and their decimal approximation was larger than mine.. so yeah i don't know where i went wrong.. the three terms (after taking the limit) came out to be e*e^3*(1/2)

OpenStudy (dan815):

oh you want to evaluate this?

OpenStudy (anonymous):

yes @dan815 and you did @TheRealMeeeee i just went about solving for it slightly differently sorry!

OpenStudy (therealmeeeee):

oh

OpenStudy (anonymous):

probably did it wrong.. it got kind of messy with everything raised to the n.. i wanted to perform the root test and then the ratio test, but i'm pretty sure that's not legal in math

OpenStudy (therealmeeeee):

well dan should get the medal dont give me one give him one

OpenStudy (dan815):

show me your work

OpenStudy (therealmeeeee):

brb dan

OpenStudy (anonymous):

ok, i started with 1. |((n+1)/(2n+7))^(n+1)| divided by a_n 2. I combined like terms (as best as i could): | ((n+1)^n)(n+1)(2n+1)^n | divided by | ((2n+7)^n)((2n+7)(n^n) | 3. simplified so i could get the limit at n-> inf | (((n+1)/n)^n)(((2n+1)/(2n+7))^(n+1))((n+1)/(2n+7))

OpenStudy (anonymous):

wait, that's the comparison test though..?

OpenStudy (anonymous):

and that converges to...0

OpenStudy (dan815):

Well it doesnt converge to 0 but ya it does converge

OpenStudy (anonymous):

what would it converge to then? i thought the lim(n->inf) (1/2n)^n is the same as (1/inf)^(inf) = 0^inf = 0

OpenStudy (dan815):

the sum wont converge

OpenStudy (anonymous):

also, does the work i showed you make sense or is it unclear? writing in equation is too long

OpenStudy (therealmeeeee):

bk bro

OpenStudy (dan815):

yep looks fine

OpenStudy (therealmeeeee):

so rember always ask the smart crew me and @dan815

OpenStudy (anonymous):

i didn't even know i could ask specific people questions.. i just post these and let any random person ask haha.. i will definitely keep you guys posted in the future. im sure youlove calc 2

OpenStudy (therealmeeeee):

lol good

OpenStudy (dan815):